OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#14037 — Optical loop in Lille
Scheduled Maintenance Report for Web Cloud
Completed
A fibre is cut on the LQN59, LZI59, ASC59, TRI59 and HEL59 lopp, on the LQN59<>LZI59 section and only in the LZI59 sense.

We've opened a ticket with the fibre operator.

Update(s):

Date: 2015-07-22 09:11:28 UTC
The problem, which came from the jumper, is now fixed.
Posted Jul 16, 2015 - 08:12 UTC